# Cross-lingual Contextualized Phrase Retrieval

### v1

@@ 6.2 Cross-lingual Contextualized Phrase Retrieval
As outlined in Section 3, given a query phrase and its context in source language, the objective of this task is to identify the most relevant cross-lingual phrase from a large-scale index in target language, emphasizing the importance of relevance not only for the query and retrieved phrases themselves but also for their respective contexts.
@@
We build the test set based on the human annotated word alignment data Jalili Sabet et al. (2020). We process the word alignment data on three language pairs, DeEn Ghader and Monz (2017), RoEn Mihalcea and Pedersen (2003), and CsEn Mareček (2011). For the CsEn dataset, we only leverage the data under the “pcedt” split, which are data from The Wall Street Journal (WSJ). For each language pair, we instruct human annotators to identify 200 high-quality phrase pairs. The details of the human annotation process are discussed in Section A.1. These annotated source phrases serve as queries, while their aligned target phrases help construct the retrieval index. To mimic real-world conditions, where a retrieval index often contains extra data, we supplement the index with additional phrases. For the uniformity of phrase segmentation across all methods, we utilize the phrases extracted by GIZA++ in the training dataset as the extra data. More concretely, we extract 9.6 million English phrases from a sample of 1 million parallel sentences from the WMT16 training dataset. Thus, each retrieval index contains all of the human-annotated target phrases and 9.6M extra phrases.
We compare our models, CCPR-XLMR and CCPR-LaBSE, with several baselines:
XLMR & mBERT: We use the pre-trained XLMR Devlin et al. (2019) and mBERT Conneau et al. (2020) models on huggingface444https://huggingface.co/models to encode the phrase contexts. We represent the phrases by concatenating the hidden representations of their start and end tokens.
mUSE & LaBSE: Compared with previous baselines, the only difference here is that we use the v2 versoin of Multilingual Universal Sentence Encoder (mUSE) Chidambaram et al. (2019) and LaBSE Feng et al. (2022) released by SBERT555https://www.sbert.net Reimers and Gurevych (2020) to encode the contexts.
CPR-X: To evaluate the effect of context-awareness, we propose two context-independent counterparts, i.e., CPR-XLMR and CPR-LaBSE. Compared with our contextualized models, the only difference is that the CPR-X methods use cross-lingual phrase pairs in semantically different contexts as positive examples, similar to the setup in Zheng et al. (2022).
For each of the models, we build an independent index for retrieval. Each of the index includes the hidden representations of human-annotated target phrases and 9.6M extra phrases encoded by the corresponding model. In this task, we use the top-1 accuracy to evaluate our methods and baselines.
@@
CCPR-XLMR | Tgt. of Train | 82.89 | 53.66 | 62.03 | 66.19 |
CCPR-XLMR | Newscrawl | 83.18 | 54.53 | 63.40 | 67.03 |
@@ Results
Table 1 demonstrates that our models significantly outperform baselines in cross-lingual contextualized phrase retrieval, highlighting the capability of our models to accurately identify relevant cross-lingual phrases while being context-aware. In contrast, the CPR-LaBSE and CPR-XLMR, i.e., the context-independent baselines, exhibit notably poorer performance.
Furthermore, despite being trained solely on sentence-level data, LaBSE still achieves moderate performance, surpassing many baselines. This supports the finding that LaBSE implicitly learns the word alignment information, as suggested by Wang et al. (2022). However, the cross-lingual contextualized phase retrieval is difficult for baselines that are only trained on multilingual data using auto-regressive loss, such as XLMR and mBERT.
@@ 6.3 Machine Translation
The evaluation on machine translation task is to assess if the information retrieved by our models can enhance performance on downstream tasks. Rather than training a new machine translation model Vaswani et al. (2017), we propose to integrate the information to the prompt of the large language model (LLM), e.g., Llama-2 Touvron et al. (2023), following the retrieval-augmented generation (RAG) methods Lewis et al. (2020a). Thus, we can fairly evaluate the effectiveness of the retrieved information.
@@
We choose the Llama-2-7B Touvron et al. (2023) model as the backbone model for the machine translation task. We also consider two cross-lingual sentence retrieval models as baseline methods, the XLM-RoBERTa-base (XLMR) model Conneau et al. (2020) and LaBSE Feng et al. (2022). An example of the instruction of our method is shown in Figure 2, where the retrieved target phrases with their contexts are presented along with the source phrases in the input to Llama-2-7B. It is worth noting that the retrieved surrounding context of each phrase is truncated to no more than 100 characters. The instructions of other baselines are shown in Appendix B.1. For all the retrieval-based methods, we build index on the monolingual newscrawl data of 2016 666https://data.statmt.org/news-crawl/${LANG}/news.2016.${LANG}.shuffled.deduped.gz, and only use the top-1 retrieved result.
The indexing strategies diverge between the baseline methods and our approach. For sentence-level baselines, i.e., XLMR and LaBSE, the indexing is straightforward, directly using the sentences. In contrast, our method employs a learned phrase segmentation module to extract phrases from sentences for indexing. When applying our model to machine translation, we set the phrase segmentation thresholds in Equation 5 to for indexing and for querying. The rationale behind a lower indexing threshold is to populate the index with a broader array of phrases, enhancing its comprehensiveness. Conversely, a higher threshold for query documents ensures that the segmented phrases are more accurate. These hyper-parameters were tuned using the validation dataset from WMT16.
Since the output format of the Llama-base model is hard to control and the Llama-chat model, i.e., the model after instruction-tuning, may reject to translate a sentence that contains negative words, we decide to fine-tune the Llama base model using LoRA Hu et al. (2022) for 100 steps to align the translation format. We use Platypus777https://github.com/arielnlee/Platypus Lee et al. (2023) as our code base. We follow all the hyper-parameters of Platypus to fine-tune the LoRA module. The data we used for fine-tuning contain 1,800 parallel sentences randomly sampled from the WMT16 training set. For each method, we only fine-tune one model for the XEn (or EnX) direction, where X is the non-English language.
@@
Main Results
As shown in Table 2, both of our CCPR-XLMR and CCPR-LaBSE outperform the baseline methods when assisting the large language model (LLM) with the information from cross-lingual contextualized phrase retrieval. Especially on some low-resource directions, such as the TrEn, EnCs, EnFi, CCPR-LaBSE outperforms other baselines by more or around 2 BERTScore points . This indicates that the the cross-lingual phrase information is more critical for improving the performance on low resource languages, because the percentage of Cs, Fi, and Tr data in Llama-2 training are all less than Touvron et al. (2023). We also evaluate our methods using additional metrics, e.g., COMET Rei et al. (2022), as shown in Appendix B.2. The results on COMET Rei et al. (2022) are consistent with those on BERTScore Zhang et al. (2019).
Additionally, we observe a good correlation between the performance of our cross-lingual contextualized phrase retrieval task and downstream tasks, such as machine translation. This suggests that models demonstrating higher retrieval accuracy tend to achieve superior results in downstream applications. Moreover, CCPR-LaBSE consistently outperforms CCPR-XLMR, highlighting the advantage of pre-training on cross-lingual sentence retrieval for enhancing the capability of contextualized phrase retrieval.
@@
In Table 2, all retrieval-based methods construct their index using the monolingual newscrawl data from the target language, which is approximately six times the size of the bilingual training data. It highlights a key benefit of our cross-lingual phrase retrieval: the ability to utilize extensive monolingual resources to build the index Cai et al. (2021), moving beyond the constraints of solely relying on bilingual data for indexing Bapna and Firat (2019); Zhang et al. (2018); Gu et al. (2018). Therefore, in Table 3, we evaluate the differences of indexing on monolingual data and the target side of the bilingual data. We observe that exploring the vast monolingual data leads to significantly better performance for both of the CCPR-XMLR and CCPR-LaBSE models.
In Table 4, we also analyze the effect of the learned phrase segmentation module (PSM). It demonstrates that PSM outperforms the -gram method by 1.8 BERTScore points in average. Although the -gram is widely used for phrase segmentation in previous works Bapna and Firat (2019); Lee et al. (2021a, b); Lan et al. (2023), our experiments demonstrate that a learned phase segmentation is more suitable for augmenting cross-lingual tasks. The explanation for this observation was discussed in Section 5.3.
@@ 7 Conclusion & Future Works
This paper presents a new formulation of dense retrieval, cross-lingual contextualized phrase retrieval, which aims to resolve the polysemy of general-type phrases by considering the context information. However, the primary issue is the lack of training data, i.e., cross-lingual phrase pairs of general type and with contexts. To address this issue, we propose to leverage the word alignment model to automatically induce word alignment information from parallel sentences and then extract cross-lingual phrase pairs for training, ensuring that extracted phrases are in general type and both phrases pairs and their contexts are aligned. Afterwards, we also introduce a Cross-lingual Contextualized Phrase Retriver (CCPR), which is mainly based on the contrastive learning, to learn close representations for cross-lingual phrases with similar meanings and surrounding contexts. We conduct extensive experiments on the cross-lingual contextualized phrase retrieval and machine translation tasks, which demonstrate that CCPR outperforms baselines by a large margin.
We believe that our cross-lingual contextualized phrase retrieval has big potential for augmenting cross-lingual tasks. In the future, we plan to focus on the acquisition of higher quality training data and on expanding our methodology to a broader range of cross-lingual applications.
@@ 9 Ethical Statement
10 Acknowledgement
We thank Leyang Cui, Patrick Lewis, Akari Asai, and Tianyu Gao for their invaluable contributions through discussions during the initial phases of this work.
Appendix A Cross-lingual Contextualized Phrase Retrieval
A.1 Human Annotation
@@
More concretely, we first use heuristic rules as discussed in Section 6.1 to collect an initial set of cross-lingual phrase pairs. For each phrase pair, we ask the human annotator to answer three questions:

### v2

@@ 6.2 Cross-lingual Contextualized Phrase Retrieval
As outlined in Sec. 3, given a query phrase and its context in source language, the objective of this task is to identify the most relevant cross-lingual phrase from a large-scale index in target language, considering the context information.
@@
We build the test set based on the human annotated word alignment data Jalili Sabet et al. (2020). We process the word alignment data on three language pairs, DeEn Ghader and Monz (2017), RoEn Mihalcea and Pedersen (2003), and CsEn Mareček (2011). For the CsEn dataset, we only leverage the data under the “pcedt” split, which are data from The Wall Street Journal (WSJ). For each language pair, we instruct human annotators to identify 200 high-quality phrase pairs. The details of the human annotation process are discussed in App. B.1. These annotated source phrases serve as queries, while their aligned target phrases help construct the index. To mimic real-world conditions, where an index contains extra data, we use 9.6 million English phrases sampled from the WMT16 training dataset as the extra data.
We compare our models, CCPR-XLMR and CCPR-LaBSE, with several baselines. The first two baselines XLMR Conneau et al. (2020) & mBERT Devlin et al. (2019) 444https://huggingface.co/models are models trained on multilingual data. We use them to encode the context and represent the phrases by concatenating the hidden representations of their start and end tokens. The second two baselines mUSE Chidambaram et al. (2019) & LaBSE Feng et al. (2022) are trained on cross-lingual data, and we use the model weights released by SBERT555https://www.sbert.net Reimers and Gurevych (2020) to encode the contexts. To evaluate the effect of context-awareness, we propose two context-independent counterparts, i.e., CPR-XLMR and CPR-LaBSE. Compared with our contextualized models, the only difference is that the CPR-X methods use cross-lingual phrase pairs in semantically different contexts as positive examples, similar to the setup in Zheng et al. (2022). We build an independent index for each model. In this task, we use the top-1 accuracy to evaluate our methods and baselines.
@@
| CCPR-XLMR | Train | 82.89 | 53.66 | 62.03 | 66.19 |
| CCPR-XLMR | NC | 83.18 | 54.53 | 63.40 | 67.03 |
@@ Results
Tab. 1 shows that our models significantly outperform baselines in cross-lingual contextualized phrase retrieval, highlighting the capability of our models to accurately identify relevant cross-lingual phrases while being context-aware. In contrast, the CPR-LaBSE and CPR-XLMR, i.e., the context-independent baselines inspired by Zheng et al. (2022), perform notably worse.
@@ 6.3 Machine Translation
The evaluation on MT task is to assess if the information retrieved by our models can enhance performance on downstream tasks. Rather than training a new MT model Vaswani et al. (2017), we propose to integrate the information to the prompt of the LLM, e.g., Llama-2 Touvron et al. (2023), following the fashion of retrieval-augmented generation (RAG) Lewis et al. (2020a). Thus, we can fairly evaluate the effect of the retrieved information.
@@
We choose Llama-2-7B Touvron et al. (2023) model as the backbone model for MT. We use two of our phrase retrieval models for this task, i.e., CCPR-Labse and CCPR-XLMR, as discussed in Sec. 6.1. We also consider a multilingual and a cross-lingual sentence retrieval models as baseline methods, i.e., the XLM-RoBERTa-base (XLMR) Conneau et al. (2020) and LaBSE Feng et al. (2022). The two sentence-level baselines are proposed to evaluate whether the phrase-level retrieval information is more beneficial for augmenting LLM’s cross-lingual ability.
We integrate the retrieval information into input of the LLM. An example of the prompt of our method is shown in Fig. 2, where the retrieved target phrases with their contexts are presented along with the source phrases in the prompt. The presented contexts are truncated to no more than 100 characters in our method. More details about prompts of our method and other baselines are shown in App. C.1. For our methods, when building the index, we first use the learned segmentation module (Sec. 5.1) to extract contextualized phrases from the retrieval data and then encode them into dense vectors using our models. For the XLMR and LaBSE baselines, we directly use them to encode sentences in retrieval data and build a sentence-level index. For all methods, we use the monolingual newscrawl data of 2016 666https://data.statmt.org/news-crawl/${LANG}/news.2016.${LANG}.shuffled.deduped.gz as the retrieval data. More setup details are in App. C.
Main Results
As shown in Tab. 2, both of our CCPR-XLMR and CCPR-LaBSE outperform the baseline methods when assisting the LLM with the information from cross-lingual contextualized phrase retrieval. Especially on some low-resource directions, such as the TrEn, EnCs, EnFi, CCPR-LaBSE outperforms other baselines by more or around 2 BERTScore points. This indicates that the the contextualized cross-lingual phrase information is more critical for improving the performance on low resource languages, because the percentage of Cs, Fi, and Tr data in Llama-2 training are all less than Touvron et al. (2023). Consistent results on COMET Rei et al. (2022) are shown in App. C.4.
In addition, we find a good correlation between the performance of our cross-lingual contextualized phrase retrieval task and the MT task, showing that higher retrieval accuracy leads to superior results in downstream applications.
@@
In Tab. 2, all retrieval-based methods construct their index using the monolingual newscrawl data, which is approximately six times the size of the bilingual training data. It highlights a key benefit of our cross-lingual phrase retrieval: the ability to utilize extensive monolingual resources to build the index Cai et al. (2021). Therefore, in Tab. 3, we evaluate the differences of indexing on monolingual data and the bilingual data. We observe that exploring the vast monolingual data leads to significantly better performance. In addition, although the -gram is widely used for phrase segmentation in previous works Bapna and Firat (2019); Lee et al. (2021a, b); Lan et al. (2023), our experiments demonstrate that a learned phase segmentation is more suitable for augmenting cross-lingual tasks, as shown in Tab. 4.
@@ 7 Conclusion & Future Works
This paper introduces a novel approach to dense retrieval, i.e., cross-lingual contextualized phrase retrieval, focusing on resolving phrase polysemy by utilizing contextual information. The main challenge identified is the scarcity of training data, specifically cross-lingual phrase pairs with context. To overcome this, we use a word alignment model to derive such phrase pairs from parallel sentences. We then present the Cross-lingual Contextualized Phrase Retriever (CCPR), which employs contrastive learning to effectively capture similar meanings and contexts of cross-lingual phrases. Our extensive testing across retrieval and machine translation tasks shows that CCPR significantly outperforms existing baselines.
@@ 9 Ethical Statement
Appendix A Implementation Details
A.1 Extraction of Cross-lingual Phrase Pairs
Because of the efficiency and the satisfactory performance of GIZA++ Och and Ney (2003); Dou and Neubig (2021), we first use the GIZA++ software777https://github.com/moses-smt/mgiza, i.e., IBM-4 model Brown et al. (1993), to induce the word alignment for each pair of parallel sentences, and then extract the cross-lingual phrase pairs. Notably, we did not induce the phrase table as in Phrase-based Statistical Machine Translation (PBSMT) Koehn et al. (2003). Instead, in our setting, we extract the cross-lingual consecutive spans that are aligned in a pair of parallel sentence as cross-lingual phrase pairs. Each of the cross-lingual phrase pairs is associated with the surrounding contexts. We filter out the phrases which begin or end with words whose frequency is more than 30k in the training dataset. We also filter out the phrases that contain only numbers and punctuations.
A.2 Training of CCPR
Each model is trained on the mixture of all language pairs of WMT16. We train our model on 8 V100 GPUs for 20K steps, where the learning rate is 5e-5, dropout rate is 0.2, batch size on each device is 64. The for the phrase segmentation loss in Eq. 5 is . Those hyper-parameters are tuned on the validation dataset.
A.3 Inference Latency
In our experiments of cross-lingual contextualized phrase retrieval, retrieving the top-32 nearest neighbors for 1000 source phrases took only 0.04 seconds on 4 V100 GPUs, thanks to FAISS’s high parallelism on GPUs Douze et al. (2024). We believe this retrieval latency is adequate for most real-time retrieval scenarios.
Appendix B Cross-lingual Contextualized Phrase Retrieval
B.1 Human Annotation
@@
More concretely, we first use heuristic rules as discussed in Sec. 6.1 to collect an initial set of cross-lingual phrase pairs. For each phrase pair, we ask the human annotator to answer three questions:
@@
Appendix C Machine Translation
C.1 Translation Instructions
C.2 Building Index
| EnDe | EnCs | EnRo | |
| CoT | 55.20 | 47.02 | 80.22 |
| CCPR-LaBSE | 64.29 | 54.93 | 83.36 |
The indexing strategies diverge between the baseline methods and our approach. For sentence-level baselines, i.e., XLMR and LaBSE, the indexing is straightforward, directly using the sentences. In contrast, our method employs a learned phrase segmentation module to extract phrases from sentences for indexing. When applying our model to MT, we set the phrase segmentation thresholds in Eq. 5 to for indexing and for querying. The rationale behind a lower indexing threshold is to populate the index with a broader array of phrases, enhancing its comprehensiveness. Conversely, a higher threshold for query documents ensures that the segmented phrases are more accurate. These hyper-parameters were tuned using the validation dataset from WMT16.
C.3 Format Alignment
Since the output format of the Llama-base model is hard to control and the Llama-chat model, i.e., the model after instruction-tuning, may reject to translate a sentence that contains negative words, we decide to fine-tune the Llama base model using LoRA Hu et al. (2022) for 100 steps to align the translation format. We use Platypus888https://github.com/arielnlee/Platypus Lee et al. (2023) as our code base. We follow all the hyper-parameters of Platypus to fine-tune the LoRA module. The data we used for fine-tuning contain 1,800 parallel sentences randomly sampled from the WMT16 training set. For each method, we only fine-tune one model for the XEn (or EnX) direction, where X is the non-English language.
C.4 More Results
More results on evaluated by COMET Rei et al. (2020) are shown in Tab. 5, which is consistent with the results evaluated by BERTScore Zhang et al. (2019).
Additional, to evaluate that the effectiveness of the retrieved information, we also compare with a Chain-of-Though (COT) approach Wei et al. (2022). As shown in Figure 8, simply translating phrases before sentence translation does not achieve the expected performance gain and is slightly worse than the baseline approach. In contrast, our approach outperforms this method by a large margin. One hypothesis is that our approach is able to leverage external knowledge, which is more helpful for the translation task.